Lupo Corporation uses a job-order costing system with a single plantwide predetermined overhead rate based on machine-hours. The company based its predetermined overhead rate for the current year on the following data:

Total machine-hours 33,000
Total fixed manufacturing overhead cost $ 660,000
Variable manufacturing overhead per machine-hour $ 6

Recently, Job T687 was completed with the following characteristics:

Number of units in the job 10
Total machine-hours 40
Direct materials $ 715
Direct labor cost $ 1,430

The amount of overhead applied to Job T687 is closest to: (Round your intermediate calculations to 2 decimal places.)

Multiple Choice

  • $208.00

  • $1,040.00

  • $1,167.40

  • $660.00
Answer #1

Total variable overhead estimated=($6*33000 machine hours)=$198000

Hence total overhead estimated=Fixed overhead+Variable overhead

=(660,000+198000)=$858,000

Hence predetermined overhead rate=Total overhead estimated/Total machine hours

=(858000/33000)=$26 per machine hour

Hence overhead applied=predetermined overhead rate*Total machine hours for the job

=(26*40 hours)

=$1040